OpenStudy (anonymous):

sharon slegde invested $20000 part at 5% and part at 7% simple interest for a period of 1 yr. How much was invested in each account if the interest earned in the 7% account was $440 greater that the interest earned in the 4% account?

OpenStudy (gw2011):

I got the following answers: $8000 was invested at 5% and $12,000 was invested at 7%. This is how I got these answers: Let x=amount invested at 5% 0.07(20,000-x)=440+0.05x 1400-0.07x=440+0.05x 1400-440=0.07x+0.05x 960=0.12x Rewriting this equation to get: 0.12x=960 x=$8000 invested at 5% $20,000-8000=$12,000 invested at 7%